# Short remarks on shallow unitary circuits

### v3

@@ Abstract
(iii) We show that any translation-invariant quantum cellular automaton in $D$-dimensional lattice of volume $V$ can be implemented using only $O(V)$ local gates in a staircase fashion.
@@ 2 Shallow circuit learning
For the last step, we recall a simple strategy that appeared in [12] where it was necessary to estimate $\operatorname{{Tr}}(UXU^{\dagger}Y)$ for various local (Pauli) operators $X$ and $Y$ given query access to $U$. Take any density matrix $\rho\succeq 0$ that, as an operator, is supported on some small set of qudits. This mixed state $\rho$ can be prepared by some local quantum channel on its support and some other channel that prepares all the qudits elsewhere in a random product state. The conjugate $U\rho U^{\dagger}$ is a density operator supported on a slightly bigger set of qudits, and usual state tomography can determine $U\rho U^{\dagger}$ whose complexity does not depend on the system size but only on the depth of UU.777 If $U$ is a shallow circuit consisting of gates from a discrete set (e.g., Clifford ++ T), then the complexity depends only on the depth. If the gates comprising $U$ is from, say, ${\mathsf{SU}}(4)$, then $U\rho U^{\dagger}$ can only be determined up to some error with some high confidence, where the accuracy and confidence must be boosted for various global requirements, and system size dependence may become necessary. If the tomography has $\sim\epsilon^{-2}\log(1/\delta)$ sample complexity for accuracy $\epsilon$ in some state distance and confidence $1-\delta$, then one would want to consume $\sim n^{2}\epsilon^{-2}\log(n/\delta)$ samples of $U\rho U^{\dagger}$ to ensure that the global accuracy of $U\otimes U^{\dagger}$ is at most $\epsilon$ with confidence $1-\delta$. We can use this to learn $(U\otimes{\mathbf{1}}){\mathsf{S}}_{i,i^{\prime}}(U^{\dagger}\otimes{\mathbf{1}})$ even though the unitary ${\mathsf{S}}_{i,i^{\prime}}$ is not a density operator. Any operator $O=\frac{O+O^{\dagger}}{2}+{\mathbf{i}}\frac{O-O^{\dagger}}{2{\mathbf{i}}}$ is a linear combination of two hermitian operators, and any hermitian operator $H$ is a linear combination of two normalized density operators:
@@ 3 Gate complexity of QCA
As discussed above, a nontrivial QCA is a quantum circuit whose spread is small but depth is large. For example, a shift QCA in one dimension can be implemented by a staircase circuit of swap gates on any circle, where the spread is $1$, but the depth is the circumference of the circle. Though it is a deep circuit, the gate count is only the number of sites. Such a circuit has been considered under the name of sequential quantum circuits [2]. We show that this is basically always the case: for any translation invariant [^10] We believe that the assumption of translation invariance is merely a technical convenience, but we do not know of a simple argument that would cover cases without translation invariance. QCA $\alpha$ of spread $r$ on a $D$-dimensional system of $n$ qudits ${\mathbb{C}}^{p}$ with periodic boundary conditions, there exists a unitary circuit $U$ consisting of $O(n)$ gates such that $\alpha(x)=UxU^{\dagger}$ for all operator $x$. Hence, in terms of gate count, every QCA is similar to a constant depth quantum circuit.
Our proof is inductive in the spatial dimension $D$ where the base case is D=1D=1.111111We could take $D=0$ as the base case, but it would be more instructive to start with $D=1$.
@@
We have converted a QCA $\alpha$ into a constant number (independent of $n$) of staircase unitary circuits, so that every qubit goes through a constant number of local unitary gates. This means that the scope of sequential unitary circuits [2] includes all translation-invariant QCA of constant spread. We did not answer whether the staircase circuit could be arranged so that all the gates were contained in a single hyperplane in spacetime of small thickness. The final circuit from our inductive proof consists of a constant number of hyperplanes in spacetime of constant thickness that contain all the local gates.

### v4

@@ Abstract
(iii) We show that any translation-invariant quantum cellular automaton in $D$-dimensional lattice of volume $V$ can be implemented using only $O(V)$ local gates in a staircase fashion using invertible subalgebra pumping.
@@ 2 Shallow circuit learning
For the last step, we recall a simple strategy that appeared in [12] where it was necessary to estimate $\operatorname{{Tr}}(UXU^{\dagger}Y)$ for various local (Pauli) operators $X$ and $Y$ given query access to $U$. Take any density matrix $\rho\succeq 0$ that, as an operator, is supported on some small set of qudits. This mixed state $\rho$ can be prepared by some local quantum channel on its support and some other channel that prepares all the qudits elsewhere in a random product state. The conjugate $U\rho U^{\dagger}$ is a density operator supported on a slightly bigger set of qudits, and usual state tomography can determine $U\rho U^{\dagger}$ whose complexity does not depend on the system size but only on the depth of UU. [^7] If $U$ is a shallow circuit consisting of gates from a discrete set (e.g., Clifford ++ T), then the complexity depends only on the depth. If the gates comprising $U$ is from, say, ${\mathsf{SU}}(4)$, then $U\rho U^{\dagger}$ can only be determined up to some error with some high confidence, where the accuracy and confidence must be boosted for various global requirements, and system size dependence may become necessary. If the tomography has $\sim\epsilon^{-2}\log(1/\delta)$ sample complexity for accuracy $\epsilon$ in some state distance and confidence $1-\delta$, then one would want to consume $\sim n^{2}\epsilon^{-2}\log(n/\delta)$ samples of $U\rho U^{\dagger}$ to ensure that the global accuracy of $U\otimes U^{\dagger}$ is at most $\epsilon$ with confidence $1-\delta$. We can use this to learn $(U\otimes{\mathbf{1}}){\mathsf{S}}_{i,i^{\prime}}(U^{\dagger}\otimes{\mathbf{1}})$ even though the unitary ${\mathsf{S}}_{i,i^{\prime}}$ is not a density operator. Any operator $O=\frac{O+O^{\dagger}}{2}+{\mathbf{i}}\frac{O-O^{\dagger}}{2{\mathbf{i}}}$ is a linear combination of two hermitian operators, and any hermitian operator $H$ is a linear combination of two normalized density operators:
@@ 3 Gate complexity of QCA
As discussed above, a nontrivial QCA is a quantum circuit whose spread is small but depth is large. For example, a shift QCA in one dimension can be implemented by a staircase circuit of swap gates on any circle, where the spread is $1$, but the depth is the circumference of the circle. Though it is a deep circuit, the gate count is only the number of sites. Such a circuit has been considered under the name of sequential quantum circuits [2]. It is shown in [2] using a characterization of QCA [22] by generalized Margolus partitioning that for any translation invariant [^10] We believe that the assumption of translation invariance is merely a technical convenience, but we do not know of a simple enough argument that would cover cases without translation invariance. QCA $\alpha$ of spread $r$ on a $D$-dimensional system of $n$ qudits ${\mathbb{C}}^{p}$ with periodic boundary conditions, there exists a unitary circuit $U$ consisting of $O(n)$ gates such that $\alpha(x)=UxU^{\dagger}$ for all operator $x$. Hence, in terms of gate count, every QCA is similar to a constant depth quantum circuit.
Here, we present another proof of the theorem of [2] by invertible subalgebra pumping. Our proof is inductive in the spatial dimension $D$ where the base case is D=1D=1.111111We could take $D=0$ as the base case, but it would be more instructive to start with $D=1$.
@@
We have converted a QCA $\alpha$ into a constant number (independent of $n$) of staircase unitary circuits, so that every qubit goes through a constant number of local unitary gates. We did not answer whether the staircase circuit could be arranged so that all the gates were contained in a single hyperplane in spacetime of small thickness. The final circuit from our inductive proof consists of a constant number of hyperplanes in spacetime of constant thickness that contain all the local gates.